OpenStudy (anonymous):

Find the solution to: x=y=7 2x-4y=8 and 5x=y=9 6x=10y=2

OpenStudy (hba):

There are two methods to solve Linear equation in two variable,and,are described as follows: 1) Substitution 2) Elimination

OpenStudy (hba):

Which method would you prefer @ChelseaChiodos ?

OpenStudy (hba):

@ChelseaChiodos Looks like you are busy :/

OpenStudy (anonymous):

equations don't make sense

OpenStudy (hba):

@ChelseaChiodos Did you try solving them up ?

OpenStudy (anonymous):

Yes, but it's confusing. They are systems of equations.

OpenStudy (anonymous):

no they just don't make sense, you have 2 equal signs in one line.

OpenStudy (hba):

I guess you have not posted the equations correctly, Therfore, i would like to see you correct the equations first.

OpenStudy (anonymous):

sorry, it's a +

OpenStudy (hba):

x+y=7 2x-4y=8 Right ?

OpenStudy (anonymous):

x+y=7 2x-4y=8 and 5x+y=9 6x+10y=2

OpenStudy (hba):

Lets see the first ques x+y=7 2x-4y=8 You can actually multiply the first equation by 2 and then subtract the two equation which will eliminate x Go on try it :)

OpenStudy (anonymous):

Okay, so 2x+2y=14 - 2x-4y=8 ?

OpenStudy (anonymous):

That would eliminate the x's and you'd be left with... -2y=22

OpenStudy (anonymous):

That makes y=-11

OpenStudy (hba):

So now you can put this -11 in eq 1 and find x ?

OpenStudy (anonymous):

(18,-11)

OpenStudy (hba):

You can check your answers by putting these values in eq 1 and 2 and if L.H.S=R.H.S then your solution is correct :)

OpenStudy (anonymous):

Thank you! :) Now, what about the 2nd problem?

OpenStudy (hba):

5x+y=9 6x+10y=2 Umm you can take value of y from equation 1 and put it in eq 2 and the solve to get x Go on show me how you do it.

OpenStudy (anonymous):

Well . . . what I did was (-10) 5x+y=9 -50x-10y=-90 6x+10y=2 Eliminated the y's and I got -44x=-88 and divided both sides by -44 I got x=22 But it didn't seem right to me :/

OpenStudy (hba):

-88/-44 = 22 ,Intelligence at its best

OpenStudy (anonymous):

So plug 22 back into the first equation... 22+y=7 Subtract 22 from both sides y=-15

OpenStudy (hba):

@ChelseaChiodos Dumb, -88/-44=2

OpenStudy (anonymous):

OOPS. sorry. I wasn't thinking.

OpenStudy (hba):

@ChelseaChiodos You need brains to do that now put this value in the 1st equation ?

OpenStudy (anonymous):

I appreciate your help, but could you not be a jerk? I'm cooking, doing PreCalc homework, and taking care of a 6 year old all at the same time. Sorry that my brain isn't entirely focused on math.

OpenStudy (anonymous):

Now, y=5.

OpenStudy (anonymous):

i believe it's called positive feedback, you're working harder now aren't you?

OpenStudy (hba):

@ChelseaChiodos I am sorry if i came of as rude but i do not agree with your accussation of calling me a jerk :) So you have done it right :) Well maths actually needs practice not attention :)

OpenStudy (anonymous):

Thanks, and I don't believe dumb is positive feedback. but okay. thank you very much, and I suppose you aren't really a jerk seems how you've helped me :P So, I apologize.

OpenStudy (hba):

@ChelseaChiodos I just came of like that so you utilize your brains,Sorry if you felt bad :)

OpenStudy (hba):

@ChelseaChiodos Your'e welcome.